Please could you help me with my homework as i am really stuck! Thanks! Here are the questions-
7x +1 = 4x +19
6x -4 = 3x +8
9x -3 = 7x +23
8x- 11= 2x +25
5x +1 = x +13
Thanks

Answer by pwac(253)

I'll do first two you try rest
7x+1=4x+19 get x-terms on left,numbers on right
7x-4x=19-1 notice sign changes when side changes,do sum
3x=18 divide both sides by 3 to get x
x=6.
6x-4=3x+8
6x-3x=8+4
3x=12
x=4.
The rest follow same method.Good Luck
Pete